Shift Summary for Eve shift, 7/Sept/99

Miguel Barrio (mbarrio@hep.uchicago.edu)
Wed, 8 Sep 1999 00:13:18 -0500

Shift Summary for Eve shift, 7/Sept/99

On Shift: BobT, Miguel Barrio

Hours of beam available at all intensities: 3.5
Hours of beam available at nominal intensity: 3.5
Hours of E832 data (special runs not included): 3.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
TeV downtime from TeV and main injector power supply problems

========================================================================
Runs ended this shift:
Run 14457 - 10.3M events - BEAM_832
Run 14458 - 6.7M events - BEAM_832

Tapes brought to FCC this shift:
QK*402-404, QKB141

========================================================================
Accesses during shift:
none

========================================================================
Overview of shift:
E832 data taking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15:45 Took tapes to Feynman (QK*402-404, QKB141).

14457 16:00 [Ed] Stopped run to install new CSQ6 constants.
New constants valid for runs [14458-99999].
References for PMT gain drift updated.

14458 16:10 Started new run.

~16:45 Beam comes back.

41 17:03 We get a DYC-FULL in the laser crate which hangs
the trigger. We stop the run and Kfrend, which
solves the problem (no need to use our
sophisticated remote powa-cycling switch).

17:19 Beam goes down, the monitor TV advertises
"TeV quench at F1 and F3"

17:47 TRD6 HV complains again: sweep current draw is
at 18uA. Almost paged Valeri, but current draw
went down to zero so we didn't.

17:51 Beam comes back.

18:46 Same DYC-FULL busy as above, again Kfrend comes
to our rescue.

20:20 Checked reference plots.

20:45 Beam goes down, some switchyard problem with a
CAMAC crate. Our SLOW-DAQ indicates that we are
not getting SEM info, which might be related to
the CAMAC crate problem MCR is having. We
checked and ktsad.exe is indeed running on
atev2.

21:05 No triggers. Kfrend-ing doesn't help.

21:10 MCR calls: Main Injector power failure, they
have no estimate for down time.

21:16 We stop run 14458, it ends ABnormally. The trigger
is hung up. Reloaded busy logic to no avail.
Various busies show up in CR53, finally DDD1->4
vetos come on. Ughhh...

We dart-fresh-start to recover, other problems
come up:
o CsI 1548 doesn't echo back correctly;
o SEM info still missing.

We are not getting RF. MCR knows about this but
they are in shambles, so they couldn't care
less.

22:20 BobT is concerned about the calorimeter not
getting RF so we took a 50MHz oscillator and we
are sending "fake" RF downstairs.

22:30 Decided to change the way we feed RF such that
all systems get it.

14459 22:45 Garbage run, nothing worked right.

14460 22:52 We are trying to initialise (notice the brit
spelling) with our fake RF just for testing
purposes. We can actually get some cosmis
triggas but no laser illumination shows up.
Calorimeter looks fine. Waiting for beam.
Estimate time for beam is 4 hours.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift:

Put real RF back into the system (when it comes back, that is);
Check that SLOW-DAQ comes back to life;
No muon run;
Debug ``No-regenerator'' trigger config;
Contact Run Manager if significant change in schedule;